# ATPrompt: Textual Prompt Learning with Embedded Attributes

### v1

@@ Abstract
Textual-based prompt learning methods primarily employ multiple learnable soft prompts and hard class tokens in a cascading manner as text prompt inputs, aiming to align image and text (category) spaces for downstream tasks. However, current training is restricted to aligning images with predefined known categories and cannot be associated with unknown categories. In this work, we propose utilizing universal attributes as a bridge to enhance the alignment between images and unknown categories. Specifically, we introduce an Attribute-embedded Textual Prompt learning method for vision-language models, named ATPrompt. This approach expands the learning space of soft prompts from the original one-dimensional category level into the multi-dimensional attribute level by incorporating multiple universal attribute tokens into the learnable soft prompts. Through this modification, we transform the text prompt from a category-centric form to an attribute-category hybrid form. To finalize the attributes for downstream tasks, we propose a differentiable attribute search method that learns to identify representative and suitable attributes from a candidate pool summarized by a large language model. As an easy-to-use plug-in technique, ATPrompt can seamlessly replace the existing prompt format of textual-based methods, offering general improvements at a negligible computational cost. Extensive experiments on 11 datasets demonstrate the effectiveness of our method.
@@ 1 Introduction
In typical image classification tasks, current text-based methods [50, 49, 15, 16] predominantly employ the traditional approach of concatenating learnable soft prompts with hard class tokens to replace handcrafted text prompts (e.g., “a photo of a {classname}”) as inputs to the encoder. Although this text prompt demonstrates strong performance, it restricts image alignment during training to predefined known categories only, thereby preventing accurate associations with unknown categories, as shown in Fig. 1(a). Intuitively, when confronted with an unfamiliar category, humans often associate it with additional attributes (e.g., color, shape, and texture) to increase comprehensibility and clarity, rather than merely stating the object’s name. For instance, one might describe a cheetah as: “The cheetah is a cat-like animal with a small head, short yellow hair, and black spots,” or refer to an apple as “That red spherical fruit with orange stripes is an apple,” instead of using a general description such as “This is a cheetah” or “That fruit is an apple.”
Inspired by these observations, in this work, we propose to utilize attributes as a bridge to enhance the alignment between images and unknown categories. Specifically, we introduce an attribute-embedded textual prompt learning method for VLMs, named ATPrompt. This method extends the learning space of soft prompts from the original one-dimensional category level to the multi-dimensional attribute level by embedding multiple fixed universal attribute tokens into the learnable soft prompts. Guided by these attributes, soft prompts acquire not only category-specific but also attribute-related general representations during training, thereby enhancing the alignment between images and unknown categories compared to the original method, as shown in Fig. 1(b). Additionally, based on the depth at which soft prompts are applied, we propose two versions of ATPrompt: shallow and deep, which ensures compatibility with existing methods of varying depths [50, 15, 16]. To finalize these universal attributes, we introduce a differentiable attribute search method, which learns to identify suitable attributes from a candidate pool constructed by LLMs. Once the search is complete, the selected attributes are integrated into ATPrompt for targeted model training.
@@
  • We propose a simple and effective attribute-embedded textual prompt learning method that expands the learning space of soft prompts from a one-dimensional class level to the multi-dimensional attribute level.
  • For selecting embedded attributes, we design a differentiable search method that learns to select suitable attributes from a candidate pool summarized by LLMs.
  • Both shallow and deep versions of ATPrompt are introduced to achieve compatibility with existing prompt learning methods of different depths.
  • Extensive experiments demonstrate that ATPrompt can be seamlessly integrated into existing textual-based methods, yielding general improvements across 11 datasets with negligible additional computational costs.
Figure 2: Architectural comparison among vanilla CLIP [33], classic prompt learning [50], and our proposed attribute-embedded prompt learning. (a) Vanilla CLIP employs a hand-crafted text template, “a photo of a {classname}”, as input to the text encoder. (b) Classical prompt learning proposes a new text form that concatenates multiple learnable soft tokens with class tokens, replacing the manually designed hard template. (c) Our ATPrompt embeds multiple fixed attribute tokens into the set of soft tokens, transforming the original form into an attribute-class mixed form for prompt learning.
@@ 2 Related Work
Prompt Learning for VLMs. Inspired by recent advancements in NLP [21, 22], prompt learning [50, 49, 51, 15, 35, 16, 24] has garnered significant interest among vision researchers aiming to apply these techniques to VLMs [33, 12], such as CLIP. CoOp [50] is the pioneering text-based approach that introduced the concept of using a combination of soft textual tokens and a hard class token as input. Subsequent studies [49, 45, 15, 20, 16, 24] have predominantly followed this textual prompt format. However, this form constrains the soft prompts to align with images within a one-dimensional, predefined category space, limiting their applicability to unknown categories. Therefore, training based on the current text form will be more likely to overfit to known categories, diminishing their zero-shot generalization capability for unknown categories. To address this limitation, several methods have been proposed [45, 14, 51, 16, 24]. For instance, KgCoOp [45] uses hand-crafted hard prompts to regularize learnable soft prompts during training. PromptSRC [16] utilizes CLIP’s [33] original features to regularize the learning of soft prompts for image and text branches. PromptKD [24] utilizes a pre-trained strong teacher model to guide the learning of a student model with learnable prompts. Despite these advancements, none of the above methods address the inherent limitations of the format itself. In this work, we introduce the attribute-embedded textual prompt format for VLMs, which proposes to utilize attributes as a bridge to build more accurate associations between images and unknown categories.
@@
Attributes for VLMs. In practice, categories typically encompass multiple attributes. When individuals encounter an unfamiliar category, they often describe it using additional attributes to enhance the clarity of their communication, rather than merely stating its name. Inspired by this observation, numerous studies [28, 4, 46, 39, 41] have begun to leverage attributes to support their objectives. VCD [28] was the pioneering work to propose the use of LLMs to decompose class names into multiple intra-class attributes (e.g., beak and tail for birds) for classification. ArGue [39] breaks down extensive descriptions into multiple intra-class attributes and trains soft prompts to align with the encoded features of these attributes. AAPL [17] introduces a meta-network to extract visual attribute features based on encoded image features, facilitating image-text alignment. Most previous studies have concentrated on utilizing intra-class attributes to improve model performance by providing supplementary attribute information. However, when confronted with an unknown class, it becomes necessary to reacquire the attributes of the new class—a process that is both complex and costly. In this work, we focus on using LLMs to mine universal inter-class attributes (e.g., color, shape, texture) for prompt learning. Instead of taking attributes as the learning objective like previous work, we propose to embed universal attributes into soft prompts, transforming the existing class-centric form [50] into a hybrid attribute-class learnable textual prompt form. Our approach can be seamlessly integrated into existing textual-based methods, enhancing their performance without incurring additional computational costs.
@@ 3 Method
Prompt learning [50, 49, 51, 15, 16] aims to enhance the generalization ability of pre-trained VLMs like CLIP on downstream tasks by training inserted learnable soft prompts. Existing textual-based methods all follow the classic prompt paradigm, concatenating soft prompt tokens and hard class tokens as the input to the encoder, as shown in Fig. 2(b). In this paper, we propose a simple and effective textual prompt learning method, named ATPrompt, which embeds multiple fixed universal attribute tokens into the original soft prompts, as shown in Fig. 2(c). Guided by these attributes, soft prompts can learn not only category-specific but also attribute-related general representations through training. When encountering unknown categories, these learned prompts can promote better image-text alignment. To determine these universal attributes, we introduce an automated pipeline consisting of multiple steps. First, we utilize LLM to summarize the attribute pool for the current downstream category. Then a differentiable attribute search method is introduced to find attributes in the pool that best suit our attribute-embedded prompt forms. Once the attributes are determined, we incorporate them into our ATPrompt for targeted model fine-tuning.
In the following, we introduce the background of VLMs and prompt learning in Sec. 3.1. Then we introduce our attribute-embedded prompt learning method in Sec. 3.2 and the attribute search method in Sec. 3.3.
Figure 3: An illustration of the computation process for shallow and deep versions. Take two attributes $[\mathrm{A}]$ and $[\mathrm{B}]$ as examples. (a) The shallow version concatenates hard attribute tokens, soft prompt tokens, and class tokens and inputs them into the encoder for calculation. (b) The deep version uses the same input but discards the class-related soft prompt tokens after calculating the self-attention and introduces them again before the next layer. These two forms can be compatible with existing methods of varying prompt depths, including input-level ones like CoOp [50], CoCoOp [49] and depth-level ones like MaPLe [15] and PromptSRC [16].
@@ 3.1 Preliminary
where M represents the soft prompt length. For simplicity, we omit the prefix and suffix tokens in the input.
@@
Figure 4: An overview of our attribute search pipeline. (a) We first query the LLM step by step to obtain multiple independent attributes, subsequently combining them to create an attribute pool for subsequent search methods. (b) Each colored path represents the forward computation process for a candidate in the pool. For attribute search, we propose the use of an alternating algorithm to jointly optimize soft tokens and path weights. After training, the attribute with the highest weight is selected for final use.
@@ 3.2 Learning soft prompts with universal attributes
Shallow Version. We begin by introducing the shallow version, in which hard attribute tokens are incorporated solely at the input level, as illustrated in Fig. 3(a). Consider two universal attributes, $\mathrm{A}$ and $\mathrm{B}$. According to Eqn. (2), the shallow-level text prompt $P_{T}$ provided to the text encoder can be expressed as follows:
$P_{T}=[T_{a_{1}}]...[T_{a_{m}}][\mathrm{A}][T_{b_{1}}]...[T_{b_{m}}][\mathrm{B}][T_{1}]...[T_{M}][\mathrm{CLS}],$ (4)
where $a_{m}$, $b_{m}$, and ${M}$ are hyperparameters specifying the length of soft tokens for attributes A, B, and the class. In our method, we set these parameters to be the same by default.
Other than placing the hard class token at the end of the text prompt, we can also place it in the middle or front like:
$P_{mid}=[T_{a_{1}}]...[T_{a_{m}}][\mathrm{A}][T_{1}]...[T_{M}][\mathrm{CLS}][T_{b_{1}}]...[T_{b_{m}}][\mathrm{B}],$ (5)
$P_{front}=[T_{1}]...[T_{M}][\mathrm{CLS}][T_{a_{1}}]...[T_{a_{m}}][\mathrm{A}][T_{b_{1}}]...[T_{b_{m}}][\mathrm{B}].$ (6)
In Tab. 4 we verify the performance of each position and select the best-performing end position as our default form.
@@ 3.3 Attribute search
To finalize the attributes for ATPrompt, two key aspects must be considered: selecting the appropriate content and determining the necessary quantity. A straightforward approach is to query the LLM directly. However, this method cannot adequately establish the optimal number of attributes for a specific downstream dataset. Additionally, without access to real data, querying solely by category name may introduce bias into the obtained attributes.
To address these challenges, we propose an automated pipeline that can select the appropriate attribute content and quantity, as illustrated in Fig. 4. Inspired by Chain-of-Thought [42, 48], we divide the entire process into multiple steps to enhance the reasoning ability of LLMs. First, we employ the LLM to generate descriptive sentences for each known category, thereby enriching category-related information. Subsequently, based on these sentences, we prompt the LLM to summarize multiple independent attribute bases across these categories. Next, an attribute pool is formed by combining different bases, as shown in Fig. 4(a). For instance, if there are five individual attribute bases, this results in 31 potential combinations of varying lengths. Note that we do not consider the attribute order specifically since it will not bring obvious semantic deviations in reality, and in the following experiments, we verify that it does not affect the final performance. We treat this pool as the search space for the following search methods.
Inspired by DARTS [25], we introduce a differentiable attribute search method that learns to find representative attributes $v$ from the search space $\mathcal{V}$, as shown in Fig. 4(b). To make the search space continuous, we relax the attribute selection operation to a softmax over all possible candidates:
@@
where $\alpha_{i}$ represents the weight for attribute $v_{i}$. The task of attribute search then reduces to learning the weight vector $\alpha$ for each attribute in the candidate pool.
@@
After relaxation, our goal is to jointly learn the attribute weight $\alpha$ and soft prompt tokens $\theta$. Same as [32, 52, 25], we use the validation set performance as the reward for the current search process and optimize the weight vector with validation loss $L_{val}$. The soft prompt tokens are obtained by minimizing the training loss $L_{train}$.
Version Shallow Shallow Shallow Deep Deep
Dataset Metric (%) CoOp +ATP CoCoOp +ATP KgCoOp +ATP MaPLe +ATP PromptSRC +ATP
Avg. Base 82.69 82.68 80.47 81.69 80.73 80.96 82.28 82.90 84.26 84.30
Novel 63.22 68.04 71.69 74.54 73.60 74.45 75.14 75.63 76.10 76.45
HM 71.66 74.65 75.83 77.95 77.00 77.57 78.55 79.10 79.97 80.18
$\Delta$ (+2.99) (+2.12) (+0.57) (+0.55) (+0.21)
Base 76.47 76.27 75.98 76.43 75.83 76.69 76.66 76.94 77.60 77.69
Image Novel 67.88 70.60 70.43 70.50 69.96 69.74 70.54 70.72 70.73 70.83
Net HM 71.92 73.33 73.10 73.35 72.78 73.05 73.47 73.70 74.01 74.10
$\Delta$ (+1.41) (+0.25) (+0.27) (+0.23) (+0.09)
Base 98.00 97.95 97.96 97.96 97.72 97.87 97.74 98.15 98.10 98.23
Caltech Novel 89.81 93.63 93.81 95.27 94.39 95.09 94.36 94.94 94.03 94.91
101 HM 93.73 95.74 95.84 96.60 96.03 96.46 96.02 96.52 96.02 96.54
$\Delta$ (+2.01) (+0.76) (+0.43) (+0.50) (+0.52)
Base 93.67 94.77 95.20 95.46 94.65 94.92 95.43 95.91 95.33 95.64
Oxford Novel 95.29 96.59 97.69 97.89 97.76 97.72 97.76 97.54 97.30 97.43
Pets HM 94.47 95.67 96.43 96.66 96.18 96.30 96.58 96.72 96.30 96.53
$\Delta$ (+1.20) (+0.23) (+0.12) (+0.14) (+0.23)
Base 78.12 77.43 70.49 74.50 71.76 72.36 72.94 74.85 78.27 79.25
Stanford Novel 60.40 66.55 73.59 73.47 75.04 74.98 74.00 73.70 74.97 74.95
Cars HM 68.13 71.58 72.01 73.98 73.36 73.65 73.47 74.27 76.58 77.04
$\Delta$ (+3.45) (+1.97) (+0.29) (+0.80) (+0.46)
Base 97.60 97.44 94.87 96.52 95.00 94.05 95.92 97.66 98.07 97.82
Flowers Novel 59.67 67.52 71.75 73.59 74.73 74.61 72.46 74.47 76.50 77.02
102 HM 74.06 79.77 81.71 83.51 83.65 83.21 82.56 84.50 85.95 86.18
$\Delta$ (+5.71) (+1.80) (-0.44) (+1.94) (+0.23)
Base 88.33 88.74 90.70 90.59 90.50 90.54 90.71 90.58 90.67 90.77
Food Novel 82.26 87.44 91.29 91.74 91.70 91.79 92.05 91.94 91.53 91.78
101 HM 85.19 88.09 90.99 91.16 91.09 91.16 91.38 91.25 91.10 91.27
$\Delta$ (+2.90) (+0.17) (+0.07) (-0.13) (+0.17)
Base 40.44 40.38 33.41 37.30 36.21 35.59 37.44 37.61 42.73 42.47
FGVC Novel 22.30 27.22 23.71 33.15 33.55 33.41 35.61 36.15 37.87 37.01
Aircraft HM 28.75 32.52 27.74 35.10 34.83 34.47 36.50 36.87 40.15 39.55
$\Delta$ (+3.77) (+7.36) (-0.35) (+0.37) (-0.60)
Base 80.60 80.84 79.74 80.50 80.29 80.09 80.82 80.96 82.67 82.73
SUN Novel 65.89 68.64 76.86 76.86 76.53 77.22 78.70 78.11 78.47 78.64
397 HM 72.51 74.24 78.27 78.64 78.36 78.63 79.75 79.51 80.52 80.63
$\Delta$ (+1.73) (+0.37) (+0.27) (-0.24) (+0.11)
DTD Base 79.44 80.83 77.01 78.63 77.55 78.30 80.36 80.33 83.37 83.22
Novel 41.18 45.49 56.00 56.89 54.99 55.32 59.18 57.85 62.97 62.68
HM 54.24 58.22 64.85 66.02 64.35 64.83 68.16 67.26 71.75 71.50
$\Delta$ (+3.98) (+1.17) (+0.48) (-0.90) (-0.25)
Base 92.19 90.34 87.49 87.95 85.64 88.27 94.07 94.84 92.90 92.29
Euro Novel 54.74 59.79 60.04 74.15 64.34 71.27 73.23 77.59 73.90 76.42
SAT HM 68.69 71.96 71.21 80.46 73.48 78.86 82.35 85.35 82.32 83.61
$\Delta$ (+3.27) (+9.25) (+5.38) (+3.00) (+1.29)
Base 84.69 84.49 82.33 82.74 82.89 81.84 83.00 84.08 87.10 87.15
UCF Novel 56.05 64.96 73.45 76.40 76.67 77.84 78.66 78.88 78.89 79.23
101 HM 67.46 73.45 77.64 79.44 79.65 79.79 80.77 81.40 82.74 83.00
$\Delta$ (+5.99) (+1.80) (+0.14) (+0.63) (+0.26)

### v2

# Advancing Textual Prompt Learning with Anchored Attributes

@@ Abstract
Textual-based prompt learning methods primarily employ multiple learnable soft prompts and hard class tokens in a cascading manner as text inputs, aiming to align image and text (category) spaces for downstream tasks. However, current training is restricted to aligning images with predefined known categories and cannot be associated with unknown categories. In this work, we propose utilizing universal attributes as a bridge to enhance the alignment between images and unknown categories. Specifically, we introduce an Attribute-anchored Textual Prompt learning method for vision-language models, named ATPrompt. This approach expands the learning space of soft prompts from the original one-dimensional category level into the multi-dimensional attribute level by incorporating multiple attribute tokens into the learnable soft prompts. Through this modification, we transform the text prompt from a category-centric form to an attribute-category hybrid form. Additionally, we introduce a straightforward differentiable attribute search method to identify representative and suitable attributes for downstream tasks. As an easy-to-use plug-in technique, ATPrompt can seamlessly replace the existing basic prompt format in textual-based methods, providing general improvements at a negligible computational cost. Extensive experiments across 11 datasets validate the effectiveness of our method.
@@ 1 Introduction
In typical image classification tasks, current text-based methods [59, 58, 18, 19] predominantly employ the traditional approach of concatenating learnable soft prompts with hard class tokens to replace handcrafted text prompts (e.g., “a photo of a {class}”) as inputs to the encoder. Although this text prompt demonstrates strong performance, it restricts image alignment during training to predefined known categories only, thereby preventing accurate associations with unknown categories, as shown in Fig. 1(a). Intuitively, when confronted with an unfamiliar category, humans often associate it with additional attributes (e.g., color, shape or texture) to increase comprehensibility and clarity, rather than merely stating the object’s name. For instance, one might describe a cheetah as: “The cheetah is a cat-like animal with a small head, short yellow hair, and black spots.” or refer to an apple as “That red spherical fruit with orange stripes is an apple.” instead of using a general description such as “This is a cheetah.” or “That fruit is an apple.” Attributes can serve as bridges that connect unknown categories to our known knowledge.
Building on these observations, we propose a novel approach that leverages attributes as a bridge to enhance the alignment between images and unknown categories. Specifically, we introduce an attribute-anchored textual prompt learning method for VLMs, named ATPrompt. This method extends the learning space of soft prompts from the original one-dimensional category level to a multi-dimensional attribute level by integrating multiple fixed universal attribute tokens into the learnable soft prompts. Guided by these anchored attributes, soft tokens acquire not only category-specific but also attribute-related general representations during training. This results in improved alignment between images and unknown categories compared to the original method, as shown in Fig. 1(b). Additionally, based on the depth at which soft prompts are applied, we propose two versions of ATPrompt: shallow and deep, respectively, ensuring compatibility with existing methods of varying depths [59, 18, 19]. To finalize these attributes, we present a simple and effective differentiable attribute search method that learns to identify suitable attributes from a candidate pool constructed by LLMs. The search operation only needs to be performed once per task, and once completed, the selected attributes can be used by ATPrompt for model training.
@@
  • We propose an efficient attribute-anchored textual prompt learning method that expands the learning space of soft prompts from a one-dimensional class level to the multi-dimensional attribute level.
  • We introduce an effective differentiable search method to select appropriate attributes for downstream tasks.
  • Both shallow and deep versions of ATPrompt are introduced to ensure compatibility with existing prompt learning methods of varying depths.
  • Extensive experiments demonstrate that ATPrompt can be seamlessly integrated into existing textual-based methods, resulting in consistently improved performance with negligible computational costs.
Figure 2: Architectural comparison among existing methods. (a) Vanilla CLIP employs a hand-crafted text template as input to the text encoder. (b) Classical prompt learning proposes a new text form that concatenates multiple learnable soft tokens with class tokens. (c) ArGue [45] employs multiple in-class attributes mined by LLMs as supplementary information. These attributes are utilized to construct distinct text groups, which serve as learning targets, thereby regularizing the learning of soft tokens. The final prediction is achieved by ensembling all groups. (d) Our ATPrompt treats universal attributes as learning components and anchors them into existing soft prompt templates. Though this operation, we expand the learning space of soft tokens to multi-dimensional attribute levels and facilitate the alignment of images with unknown class texts.
@@ 2 Related Work
Prompt Learning for VLMs. Inspired by recent advancements in NLP [24, 25], prompt learning [59, 58, 60, 18, 40, 19, 27, 56] has garnered significant interest among vision researchers aiming to apply these techniques to VLMs [38, 15, 51], such as CLIP. CoOp [59] is the pioneering text-based approach that introduced the concept of using a combination of soft textual tokens and a hard class token as input. Subsequent studies [58, 53, 18, 23, 19, 27] have predominantly followed this textual prompt format. However, this form constrains the soft prompts to align with images within a one-dimensional, predefined category space, limiting their applicability to unknown categories. Therefore, training based on the current text form will be more likely to overfit to known categories, diminishing their zero-shot generalization capability for unknown categories. To address this limitation, several methods have been proposed [53, 17, 60, 19, 27]. For instance, KgCoOp [53] uses hand-crafted hard prompts to regularize learnable soft prompts during training. PromptSRC [19] utilizes CLIP’s [38] original features to regularize the learning of soft prompts for image and text branches. PromptKD [27] utilizes a pre-trained strong teacher model to guide the learning of a student model [13, 26, 52, 28] with learnable prompts. Despite these advancements, none of the above methods address the inherent limitations of the format itself. In this work, we introduce the attribute-anchored textual prompt format for VLMs, which proposes to utilize attributes as a bridge to build more accurate associations between images and unknown categories.
@@
Attributes for VLMs. In practice, categories typically encompass multiple attributes. When individuals encounter an unfamiliar category, they often describe it using additional attributes to enhance the clarity of their communication, rather than merely stating its name. Inspired by this observation, numerous studies [33, 4, 54, 45, 47] have begun to leverage attributes to support their objectives. VCD [33] was the pioneering work to propose the use of LLMs to decompose class names into multiple in-class attributes (e.g., beak and tail for birds) for classification. AAPL [20] introduces a meta-network to extract visual attribute features based on encoded image features, facilitating image-text alignment. TAP [7] presents a structured “Tree of Attributes” approach that leverages attribute-specific knowledge graphs to enhance VLMs. ArGue [45] leverages large language models to mine multiple in-class attributes and integrates them into soft prompts and fixed templates to create multiple text groups. It employs the original text features generated by the fixed template to regularize the learning of soft tokens, as illustrated in Fig. 2(c). Most prior studies have focused on utilizing in-class attributes to enhance model performance by providing supplementary attribute information. However, when dealing with an unknown class, reacquiring the attributes of the new class becomes necessary—a process that is both complex and costly.
In this work, we believe that universal (inter-class) attributes are more efficient and robust than the in-class attributes used in previous works. Instead of taking attributes as the learning objective, we treat it as a learning component and propose to anchor universal attributes into soft prompt templates, transforming the existing class-centric form [59] into a hybrid attribute-class learnable textual prompt form. Our approach can be seamlessly integrated into existing textual-based methods, enhancing their performance without incurring additional computational costs.
@@ 3 Method
Prompt learning [59, 58, 60, 18, 19] aims to enhance the generalization ability of pre-trained VLMs like CLIP on downstream tasks by training inserted learnable soft tokens. Existing textual-based methods all follow the classic prompt paradigm, concatenating soft prompt tokens and hard class tokens as the input to the text encoder, as shown in Fig. 2(b). In this paper, we propose a simple and effective textual prompt learning method, named ATPrompt, which anchors multiple fixed universal attribute tokens into the original soft prompts, as shown in Fig. 2(d). Guided by these attributes, soft prompts can learn not only category-specific but also attribute-related general representations through training. When encountering unknown categories, these learned attribute-related tokens can provide additional information to promote better image-text alignment. Furthermore, to identify these universal attributes, we present an automated pipeline that encompasses sequential steps. Initially, we employ LLMs to synthesize the attribute pool for the current downstream category. Subsequently, we propose a differentiable attribute search method designed to identify attributes within the pool that are most suitable for our attribute-anchored prompt forms. For every task, this search operation is conducted only once. Once the attributes are finalized, they are integrated into our ATPrompt for specialized model fine-tuning.
Figure 3: An illustration of the computation process for shallow and deep versions. Take two attributes $[\mathrm{A}]$ and $[\mathrm{B}]$ as examples. (a) The shallow version concatenates hard attribute tokens, soft prompt tokens, and class tokens and inputs them into the encoder for calculation. (b) The deep version uses the same input but discards the class-related soft prompt tokens after calculating the self-attention and introduces them again before the next layer. These two forms can be compatible with existing methods of varying prompt depths, including input-level ones like CoOp [59], CoCoOp [58] and depth-level ones like MaPLe [18].
@@ 3.1 Preliminary
where $M$ represents the length of soft tokens. For simplicity, we omit the prefix and suffix tokens in the input.
@@
Figure 4: An overview of our attribute search pipeline. (a) We first query the LLM iteratively to obtain multiple independent attributes. These are subsequently aggregated to form a pool of candidate combinations, which serves as the input for the search process. (b) The forward computation for each candidate combination is represented by a distinct colored path. To identify the optimal attributes, we employ an alternating optimization algorithm that co-optimizes the soft tokens and a corresponding path weight vector. Upon completion of training, the combination associated with the highest-weighted path is selected as the final output.
@@ 3.2 Learning soft prompts with universal attributes
Shallow Version. We begin by introducing the shallow version, in which hard attribute tokens are anchored solely at the input level, as illustrated in Fig. 3(a). Consider two universal attributes, $\mathrm{A}$ and $\mathrm{B}$. According to Eqn. (2), the shallow-level text prompt $P_{T}$ provided to the text encoder can be expressed as follows:
$P_{T}=[T_{a_{1}}]...[T_{a_{m}}][\mathrm{A}][T_{b_{1}}]...[T_{b_{m}}][\mathrm{B}][T_{1}]...[T_{M}][\mathrm{CLS}].$ (4)
where $a_{m}$ and $b_{m}$ are hyperparameters specifying the length of soft tokens for attributes A and B. In our method, we set these parameters to be the same by default.
Other than placing the hard class token at the end of the text prompt, we can also place it in the middle or front. In Tab. 5, we verify the performance of each position and select the best-performing end position as our default form.
@@ 3.3 Attribute search
Selecting the attributes involves two key considerations: their content and their quantity. While directly querying a LLM is a straightforward approach, it has notable drawbacks. This method cannot determine the optimal number of attributes for a specific downstream dataset, and querying by category name alone can introduce semantic bias. To address these issues, we propose an automated pipeline that selects the appropriate content and quantity of attributes for the current downstream task, as illustrated in Fig. 4.
Attribute Pool. Inspired by CoT [48, 57], we divide the entire process into multiple steps to enhance the reasoning ability of LLMs. First, we prompt the LLM to generate descriptive sentences for each known category, thereby enriching category-related information. Using these descriptions as context, we then prompt the LLM to summarize a set of independent attribute bases that are common across these categories. An attribute pool is then formed by creating all possible combinations of these bases, as shown in Fig. 4(a). For $N$ attribute bases, this result in a total of $w=C_{N}^{1}+C_{N}^{2}+...+C_{N}^{N}$ candidates in the pool, which constitutes our search space. Note that we do not consider the order of attributes, as permutations generally do not introduce significant semantic bias or affect the final performance, a claim we validate in our experiments.
Attribute Searching. Inspired by DARTS [30], we introduce a differentiable attribute search method that learns to find representative attributes $v$ from the search space $\mathcal{V}$, as shown in Fig. 4(b). To make the search space continuous, we relax the discrete attribute selection into a softmax-weighted sum over all $w$ possible candidates:
@@
where $\alpha_{i}$ represents the weight for attribute combination $v_{i}$. The task of attribute search is thus reduced to learning the weight vector $\alpha$ for the candidate pool.
@@
After relaxation, our goal is to jointly learn the attribute weight $\alpha$ and soft prompt tokens $\theta$. Following standard practice [37, 61, 30], we optimize the weights $\alpha$ by minimizing the validation loss $L_{val}$, while the soft tokens are learned by minimizing the training loss $L_{train}$. We employ an alternating algorithm [14, 26] to solve this bi-level optimization problem, alternating between these two subpromblems:
Method Average ImageNet Caltech101 OxfordPets
Base Novel HM Base Novel HM Base Novel HM Base Novel HM
CoOp (IJCV 22) 82.69 63.22 71.66 76.47 67.88 71.92 98.00 89.81 93.73 93.67 95.29 94.47
CoCoOp (CVPR 22) 80.47 71.69 75.83 75.98 70.43 73.10 97.96 93.81 95.84 95.20 97.69 96.43
MaPLe (CVPR 23) 82.28 75.14 78.55 76.66 70.54 73.47 97.74 94.36 96.02 95.43 97.76 96.58
PromptSRC (ICCV 23) 84.26 76.10 79.97 77.60 70.73 74.01 98.10 94.03 96.02 95.33 97.30 96.30
ArGue (CVPR 24) 83.69 78.07 80.78 76.92 72.06 74.41 98.43 95.20 96.79 95.36 97.95 96.64
DePT (CVPR 24) 83.66 71.82 77.29 77.13 70.10 73.45 98.33 94.33 96.29 94.70 97.63 96.14
CoPrompt (ICLR 24) 84.00 77.23 80.48 77.67 71.27 74.33 98.27 94.90 96.55 95.67 98.10 96.87
PromptKD (CVPR 24) 86.96 80.73 83.73 80.83 74.66 77.62 98.91 96.65 97.77 96.30 98.01 97.15
CoOp + ATPrompt 82.68 68.04 74.65 (+2.99) 76.27 70.60 73.33 97.95 93.63 95.74 94.77 96.59 95.67
CoCoOp + ATPrompt 81.69 74.54 77.95 (+2.12) 76.43 70.50 73.35 97.96 95.27 96.60 95.46 97.89 96.66
MaPLe + ATPrompt 82.98 75.76 79.21 (+0.66) 76.94 70.72 73.70 98.32 95.09 96.68 95.62 97.63 96.61
DePT + ATPrompt 83.80 73.75 78.45 (+1.16) 77.32 70.65 73.83 98.48 94.60 96.50 94.65 97.99 96.29
PromptKD + ATPrompt 87.05 81.82 84.35 (+0.62) 80.90 74.83 77.75 98.90 96.52 97.70 96.92 98.27 97.59
Method StanfordCars Flowers102 Food101 FGVCAircraft
Base Novel HM Base Novel HM Base Novel HM Base Novel HM
CoOp (IJCV 22) 78.12 60.40 68.13 97.60 59.67 74.06 88.33 82.26 85.19 40.44 22.30 28.75
CoCoOp (CVPR 22) 70.49 73.59 72.01 94.87 71.75 81.71 90.70 91.29 90.99 33.41 23.71 27.74
MaPLe (CVPR 23) 72.94 74.00 73.47 95.92 72.46 82.56 90.71 92.05 91.38 37.44 35.61 36.50
PromptSRC (ICCV 23) 78.27 74.97 76.58 98.07 76.50 85.95 90.67 91.53 91.10 42.73 37.87 40.15
ArGue (CVPR 24) 75.64 73.38 74.49 98.34 75.41 85.36 92.33 91.96 92.14 40.46 38.03 39.21
DePT (CVPR 24) 79.67 72.40 75.86 98.20 72.00 83.08 90.43 91.33 90.88 42.53 22.53 29.46
CoPrompt (ICLR 24) 76.97 74.40 75.66 97.27 76.60 85.71 90.73 92.07 91.40 40.20 39.33 39.76
PromptKD (CVPR 24) 82.80 83.37 83.13 99.42 82.62 90.24 92.43 93.68 93.05 49.12 41.81 45.17
CoOp + ATPrompt 77.43 66.55 71.58 97.44 67.52 79.77 88.74 87.44 88.09 40.38 27.22 32.52
CoCoOp + ATPrompt 74.50 73.47 73.98 96.52 73.59 83.51 90.59 91.74 91.16 37.30 33.15 35.10
MaPLe + ATPrompt 75.39 73.84 74.61 97.82 75.07 84.95 90.65 92.00 91.32 37.61 36.15 36.87
DePT + ATPrompt 79.29 73.47 76.27 98.20 73.69 84.20 90.42 91.69 91.05 43.19 33.23 37.56
PromptKD + ATPrompt 82.51 84.03 83.26 99.15 82.03 89.78 92.48 93.86 93.22 49.63 42.35 45.70
Method SUN397 DTD EuroSAT UCF101
Base Novel HM Base Novel HM Base Novel HM Base Novel HM
CoOp (IJCV 22) 80.60 65.89 72.51 79.44 41.18 54.24 92.19 54.74 68.69 84.69 56.05 67.46
CoCoOp (CVPR 22) 79.74 76.86 78.27 77.01 56.00 64.85 87.49 60.04 71.21 82.33 73.45 77.64
MaPLe (CVPR 23) 80.82 78.70 79.75 80.36 59.18 68.16 94.07 73.23 82.35 83.00 78.66 80.77
PromptSRC (ICCV 23) 82.67 78.47 80.52 83.37 62.97 71.75 92.90 73.90 82.32 87.10 78.80 82.74
ArGue (CVPR 24) 81.52 80.74 81.13 81.60 66.55 73.31 94.43 88.24 91.23 85.56 79.29 82.31
DePT (CVPR 24) 82.37 75.07 78.55 83.20 56.13 67.04 88.27 66.27 75.70 85.43 72.17 78.24
CoPrompt (ICLR 24) 82.63 80.03 81.30 83.13 64.73 72.79 94.60 78.57 85.84 86.90 79.57 83.07
PromptKD (CVPR 24) 83.69 81.54 82.60 85.84 71.37 77.94 97.54 82.08 89.14 89.71 82.27 86.10
CoOp + ATPrompt 80.84 68.64 74.24 80.83 45.49 58.22 90.34 59.79 71.96 84.49 64.96 73.45
CoCoOp + ATPrompt 80.50 76.86 78.64 78.63 56.89 66.02 87.95 74.15 80.46 82.74 76.40 79.44
MaPLe + ATPrompt 80.98 78.15 79.54 80.50 58.28 67.61 94.84 77.59 85.35 84.08 78.88 81.40
DePT + ATPrompt 82.42 76.48 79.34 82.64 56.77 67.30 89.60 69.50 78.28 85.60 73.15 78.89
PromptKD + ATPrompt 83.87 81.35 82.59 86.92 72.34 78.96 97.05 92.07 94.49 89.29 82.44 85.73